OpenStudy (mislovely187):

Find the radius and center of the circle given by the equation below. (x – 6)2 + (y + 4)2 = 7

OpenStudy (danjs):

circle at origin, radius r x^2+y^2 = r^2 circle radius r, shifted to have center at a point(h,k) \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (danjs):

what you think it is from that?

OpenStudy (mislovely187):

im confuesd :-(